How to solve:
The differences between the square of a positive number and ten times the number 11. Find the positive number.

How to solve:
The height of a parallelogram is 2cm longer than its base. Its area is 120cm2. Find the base and height of the parallelogram.

Answer by ankor@dixie-net.com(22740)

The differences between the square of a positive number and ten times the number is 11.
Find the positive number.
x^2 - 10x = 11
x^2 - 10x - 11 = 0; our old friend, the quadratic equation!
Factors to
(x-11)(x+1) = 0
The positive solution
x = 11
Check: 11^2 - 10(11) = 11
:
How to solve:
The height of a parallelogram is 2cm longer than its base.
Its area is 120cm2. Find the base and height of the parallelogram.
:
Area of a parallelogram;
b * h = 120
"height is 2cm longer than its base."
h = (b+2)
Replace h with (b+2) in the area equation
b * (b+2) = 120
b^2 + 2b = 120
b^2 + 2b - 120 = 0; what luck! Another quadratic equation!
Factors to
(b-10)(b+12) = 0
Positive solution
b = 10 cm
then
h = 10+2 = 12 cm
;
Check 10 * 12 = 120 sq/cm